# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ Abstract
A well-defined distance on the parameter space is key to evaluating estimators, ensuring consistency, and building confidence sets. While there are typically standard distances to adopt in a continuous space, this is not the case for combinatorial parameters such as graphs that represent statistical models. Existing proposals like the structural Hamming distance are defined on the graphs rather than the models they represent and can hence lead to undesirable behaviors. We propose a model-oriented framework for defining the distance between graphs that is applicable across many different graph classes. Our approach treats each graph as a statistical model and organizes the graphs in a partially ordered set based on model inclusion. This induces a neighborhood structure, from which we define the model-oriented distance as the length of a shortest path through neighbors, yielding a metric in the space of graphs. We apply this framework to both probabilistic graphical models (e.g., undirected graphs and completed partially directed acyclic graphs) and causal graphical models (e.g., directed acyclic graphs and maximally oriented partially directed acyclic graphs). We analyze the theoretical and empirical behaviors of model-oriented distances. Algorithmic tools are also developed for computing and bounding these distances.
@@ 1 Introduction
Graphs provide an intuitive framework for encoding and reasoning about the dependency structure among a set of variables. In statistics, they have become popular in two closely related but different contexts: probabilistic modeling and causal modeling. For the former context, graphs typically encode conditional independence relations among a set of (factual) random variables (Lauritzen, 1996); for certain scenarios, they can also encode generalized versions of conditional independence and other constraints (Evans, 2016; Richardson et al., 2023) — we refer to these as probabilistic graphical models. Among other applications, these graphs lead to sound and efficient algorithms for probabilistic reasoning (Lauritzen and Spiegelhalter, 1988; Lunn et al., 2000). For the latter context, a graph models a causal system by encoding the relations of its building blocks, namely counterfactual (as well as factual) random variables; as such, the graph not only describes the system’s behavior as things naturally occur, but also its behavior under interventions (Wright, 1934; Spirtes, Glymour and Scheines, 2000; Pearl, 2009) — we use the term causal graphical models to refer to graphs like this. These graphs are used to determine if a causal effect can be identified from observational data (Huang and Valtorta, 2006; Shpitser and Pearl, 2006), and if so, to also inform the best way to estimate the effect (Rotnitzky and Smucler, 2020; Henckel, Perković and Maathuis, 2022; Guo, Perković and Rotnitzky, 2023). There is also a large literature on learning probabilistic or causal graphs from data, a problem known as structure learning or causal discovery; the reader is referred to Drton and Maathuis (2017); Glymour, Zhang and Spirtes (2019) for two recent surveys.
@@
In whichever context these graphs are used, a key unresolved issue is how to define a suitable distance between them. Having a distance is essential for several tasks, including benchmarking causal discovery algorithms (Scutari, Graafland and Gutiérrez, 2019; Rios, Moffa and Kuipers, 2023), characterizing their statistical performance (Jin, Zhang and Zhang, 2012; Ke, Jin and Fan, 2012; Han et al., 2016) or finding a consensus graph among a collection of graphs through a Frechét mean (Ferguson and Meyer, 2023; Wang, Kolar and Drton, 2025). Other potential applications include constructing ball-like confidence regions for a graph and and conducting sensitivity analysis with respect to potential mistakes in the causal graph. However, defining a suitable distance is challenging for two reasons. First, there are many types of graphs, such as undirected graphs (UGs), directed acyclic graphs (DAGs), completed partially directed acyclic graphs (CPDAGs), and maximally oriented partially directed acyclic graphs (MPDAGs), each with its own edge types and rules on how edges may be configured. Second, each type of graphs rely on a set of semantics (sometimes called Markov properties) to be interpreted as probabilistic or causal models, and a good distance between two graphs should reflect how close they are in terms of the models they represent. That is, we seek a semantic rather than structural definition of the distance. Despite these challenges, several distances have been proposed in the literature (Tsamardinos, Brown and Aliferis, 2006; Peters and Bühlmann, 2015; Viinikka, Eggeling and Koivisto, 2018), which we divide into two categories — model-agnostic and task-oriented distances.
@@
Arguably, the most popular distance is the structural Hamming distance (SHD) (Tsamardinos, Brown and Aliferis, 2006). It counts the number of edges that differ between two graphs with various proposals for how to weigh different types of edge mismatches (Perrier, Imoto and Miyano, 2008). Irrespective of the concrete weighing the SHD is computationally inexpensive, flexible with respect to the graph class, and guaranteed to define a metric. It is, however, model-agnostic — it disregards the specific models the graphs represent and is solely defined on the graphs themselves. This results in undesirable behaviors. First, for a pair of graphs, the SHD provides a fixed answer regardless of how the graphs are interpreted, whether probabilistically or causally. Second, the SHD does not depend on the set of graphs in consideration — it is agnostic to constraints posed by the graph type and domain specific considerations (e.g., one may wish to consider acyclic graphs). Finally, by construction the SHD treats each edge individually. This ignores that not all edge configurations may give a valid graph in the graph class under consideration. We illustrate these issues with an example.
@@
Consider the four CPDAGs in Fig.˜1(a-d). For two of the more prominent types of SHD (see Section˜4.3), we have $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{2})=n+1$, respectively $n+2$. Similarly, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{3})=2\text{ or }4$, and for even $n$, $\mathrm{SHD}(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1\text{ or }n+2$. However, in the context of probabilistic models, $\mathcal{G}_{2}$ sits “just between” $\mathcal{G}_{1}$ and $\mathcal{G}_{3}$: $\mathcal{G}_{2}$ encodes fewer conditional dependencies than $\mathcal{G}_{1}$ and no graph encodes an intermediary collection of dependencies. The same is true for $\mathcal{G}_{3}$ with respect to $\mathcal{G}_{2}$. Logically, $\mathcal{G}_{3}$ should be closer to $\mathcal{G}_{2}$ than $\mathcal{G}_{1}$. The SHD overestimates the semantic distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, as it wrongly assumes that undirecting any edge in $\mathcal{G}_{2}$ leads to a valid CPDAG that is “in between” $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Similarly, since $\mathcal{G}_{4}$ removes far more dependencies from $\mathcal{G}_{1}$ than $\mathcal{G}_{2}$ does, we expect $\mathcal{G}_{1}$ to be closer to $\mathcal{G}_{2}$ than to $\mathcal{G}_{4}$. Our proposed model-oriented distance reflects the semantics of the graphs, and yields $d(\mathcal{G}_{1},\mathcal{G}_{2})=1$, $d(\mathcal{G}_{1},\mathcal{G}_{3})=2$, and $d(\mathcal{G}_{1},\mathcal{G}_{4})=n/2+1$. As a further illustration, we set $n=10$, and generate $100$ datasets of $1000$ samples each from a linear Gaussian structural equation model from graph $\mathcal{G}_{1}$, with unit edge weight and noise variance. We can fit a given graph $\mathcal{G}$ to the data and compute the Bayesian Information Criterion (BIC), which balances model fit and complexity. As shown in Fig.˜1(e-f), our proposed model-oriented distance better reflects fit to data: further away graphs tend to yield worse BIC scores. This indicates that using SHD to benchmark score-based causal discovery algorithms that rely on BIC, such as the well-known greedy equivalence search algorithm (Chickering, 2002), may be misleading.
@@
To address the limitation of model-agnostic distances, Peters and Bühlmann (2015) proposed an alternative, task-oriented approach, based on the idea that the distance between two graphs should be measured based on the specific task they serve. In the context of causal inference, the central task is to identify a causal effect of interest, i.e., to express the effect as a functional of the observed distribution that holds under every causal system that is compatible with a graph. Peters and Bühlmann (2015) introduced the structural intervention distance (SID) between two DAGs, which counts how many identifying formulae under one graph would be invalid if the truth lies within the other graph. This approach was also adapted and generalized by Henckel, Würtzen and Weichwald (2024) who establish that the SID is merely one of a class of closely related distances they call adjustment identification distances. The SID is, however, not a metric: it can be zero between two different graphs, is not symmetric and does not obey the triangle inequality. It is possible to define a symmetrized version of SID but we show in Section˜H.1 that this version still violates the triangle inequality. In general task-oriented distance are not constructed to be metrics, limiting their usefulness.
@@
(e) SHD distance vs BIC score
(f) Proposed model-oriented distance vs BIC score
Figure 1: (a–f) Four CPDAGs in Example˜1 and their associated BIC scores for synthetically generated data.
1.1 Our contributions
We introduce a model-oriented framework that defines distance between graphs through the models they represent. Given a class of graphs of the same type (e.g., the collection of CPDAGs over five vertices), our approach begins by interpreting a probabilistic or causal graph as a statistical model, namely a set of distributions over a common collection of random variables — factual or counterfactual — that obey the constraints encoded by the graph. This viewpoint naturally leads to the organization of graphs into a partially ordered set (poset), where the ordering is defined by model inclusion: for two graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$, we say $\mathcal{G}\preceq\mathcal{G}^{\prime}$ if and only if the model represented by $\mathcal{G}$ is contained by the model represented by $\mathcal{G}^{\prime}$. Posets are equipped with a natural neighborhood notion called covering elements. Exploiting this neighborhood notion, we define our distance as the length of a shortest path between two graphs when moving only through neighbors. By construction, our distance is a metric and applicable to many classes of graphs of varying type.
We study in detail the model-oriented distance for four graph types, listed in order of increasing complexity of their associated posets: probabilistic UGs, causal DAGs, probabilistic CPDAGs, and causal MPDAGs. For each, we study algorithms for computing or bounding the model-oriented distance. Here, a key ingredient is a graphical characterization of model containment, and whenever possible, of neighborhood. When the latter proves difficult, we study the neighborhood for a restricted class of graphs where the characterization is tractable. Further, we show that poset-theoretic properties, such as gradedness and semimodularity, can facilitate computing the distances. We also show that, for CPDAGs and MPDAGs, the model-oriented distance can differ substantially from the structural Hamming distance. For CPDAGs and MPDAGs restricted to polytrees, we also prove that the model-oriented distance admits several simplifications that make it easier to compute and interpret.
By establishing that many statistical graph classes are naturally equipped with partial orderings that can be studied using theoretical tools from order theory, we open a new avenue for investigating graphical models and their properties. Partial orders, for example, have been shown to be natural and useful structures for model selection over combinatorial objects, including graphs (Taeb, Bühlmann and Chandrasekaran, 2023). Similarly, neighborhood notions are central to score-based causal discovery methods where they used to decide how to explore the space of all graphs in a class while searching for a score-minimizing graph (Chickering, 2002).
@@ 2.1 Graphical preliminaries
We consider graphs with nodes representing random variables and edges between them representing probabilistic or causal dependencies between these variables. Basic graphical concepts are briefly described in below.
2.1.1 Node, edge, subgraph, cycle, PDAG
We write a graph over vertex set $\mathrm{V}$ with $k\geq 1$ types of edges (e.g., undirected ‘-’ and directed $\rightarrow$’) as $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$. Depending on the edge type $i$, edge set $\mathrm{E}^{i}$ is either a set of unordered tuples ($(1,2)=(2,1)$) or a set of ordered tuples ($(1,2)\neq(2,1)$). In either case, we forbid self-loops and let $\mathrm{E}^{i}\subseteq(\mathrm{V}\times\mathrm{V})\setminus\cup_{i\in\mathrm{V}}\{(i,i)\}$ for $i=1,\dots,k$. A graph $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}^{1},\dots,\mathrm{E}_{1}^{k})$ is said to be a subgraph of $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2}^{1},\dots,\mathrm{E}_{2}^{k})$ if $\mathrm{E}_{1}^{i}\subseteq\mathrm{E}_{2}^{i}$ for each edge type $i$. Given a graph $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$ and a subset of vertices $\mathrm{V}^{\prime}\subseteq\mathrm{V}$, the induced graph of $\mathrm{V}^{\prime}$ is $\mathcal{G}_{\mathrm{V}^{\prime}}:=(\mathrm{V}^{\prime},{\mathrm{E}^{1}}^{\prime},\dots,{\mathrm{E}^{k}}^{\prime})$ where ${\mathrm{E}^{i}}^{\prime}:=\mathrm{E}^{i}\cap(\mathrm{V}^{\prime}\times\mathrm{V}^{\prime})$ for $i=1,\dots,k$. A graph is simple if between every two vertices there is at most one edge (of any type). Throughout, we only consider simple graphs. A cycle is a sequence of vertices $v_{1},v_{2},\dots,v_{k}$ ($k\geq 3$) such that there is an edge between every two adjacent vertices, $v_{k}=v_{1}$ and $v_{1},\dots,v_{k-1}$ are all distinct. A directed cycle is a cycle of the form $v_{1}\rightarrow v_{2}\rightarrow\dots\rightarrow v_{k}$. In contrast, a partially directed cycle is a cycle where the edges can be directed or undirected such that (i) at least one edge is directed and (ii) all the directed edges in the cycle point to the same direction. A graph is chordal if every cycle of length 4 or more has a chord — an edge connecting two non-consecutive vertices. A partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.

### v2

@@ Abstract
A well-defined distance on the parameter space is key to evaluating estimators, ensuring consistency, and building confidence sets. While there are typically standard distances to adopt in a continuous space, this is not the case for combinatorial parameters such as graphs that represent statistical models. Defined on the graphs alone, existing proposals like the structural Hamming distance ignore the structure of the model space and can thus exhibit undesirable behaviors. We propose a model-oriented framework for defining the distance between graphs that is applicable across different graph classes. Our approach treats each graph as a statistical model and organizes the graphs in a partially ordered set based on model inclusion. This induces a neighborhood structure, from which we define the model-oriented distance as the length of a shortest path through neighbors, yielding a metric in the space of graphs. We apply this framework to probabilistic undirected graphs, causal directed acyclic graphs, probabilistic completed partially directed acyclic graphs, and causal maximally oriented partially directed acyclic graphs. We analyze theoretical and empirical behaviors of the model-oriented distance and draw comparison with existing distances. By exploiting the underlying poset structures, we develop algorithms for computing and bounding the proposed distance that scale to moderate-sized graphs.
@@ 1 Introduction
Graphs provide an intuitive framework for encoding and reasoning about the dependency structure among a set of variables. In statistics, they have become popular in two closely related but different contexts: probabilistic modeling and causal modeling. For the former context, graphs typically encode conditional independence relations among a set of (factual) random variables (Lauritzen, 1996); for certain scenarios, they can also encode generalized versions of conditional independence and other constraints (Evans, 2016; Richardson et al., 2023) — we refer to these as probabilistic graphical models. Among other applications, these graphs lead to sound and efficient algorithms for probabilistic reasoning (Lauritzen and Spiegelhalter, 1988; Lunn et al., 2000). For the latter context, a graph models a causal system by encoding the relations of its building blocks, namely counterfactual (as well as factual) random variables; as such, the graph not only describes the system’s behavior as things naturally occur, but also its behavior under interventions (Wright, 1934; Spirtes, Glymour and Scheines, 2000; Pearl, 2009) — we use the term causal graphical models to refer to graphs like this. These graphs are used to determine if a causal effect can be identified from observational data (Huang and Valtorta, 2006; Shpitser and Pearl, 2006), and if so, to also inform the best way to estimate the effect (Rotnitzky and Smucler, 2020; Henckel, Perković and Maathuis, 2022; Guo, Perković and Rotnitzky, 2023). There is also a large literature on learning probabilistic or causal graphs from data, a problem known as structure learning or causal discovery; the reader is referred to Drton and Maathuis (2017); Glymour, Zhang and Spirtes (2019) for two surveys.
@@
In whichever context these graphs are used, a key unresolved issue is how to define a suitable distance between them. Having a distance is essential for several tasks, including benchmarking causal discovery algorithms (Scutari, Graafland and Gutiérrez, 2019; Rios, Moffa and Kuipers, 2025), characterizing their statistical performance (Jin, Zhang and Zhang, 2012; Ke, Jin and Fan, 2012; Han et al., 2016) or finding a consensus graph among a collection of graphs through a Frechét mean (Ferguson and Meyer, 2023; Wang, Kolar and Drton, 2025). Other potential applications include constructing ball-like confidence regions for a graph and conducting sensitivity analyses with respect to potential mistakes in the causal graph. However, defining a suitable distance is challenging for two reasons. First, there are many types of graphs, such as undirected graphs (UGs), directed acyclic graphs (DAGs), completed partially directed acyclic graphs (CPDAGs), and maximally oriented partially directed acyclic graphs (MPDAGs), each with its own edge types and rules on how edges may be configured. Second, each type of graphs rely on a set of semantics (sometimes called Markov properties) to be interpreted as probabilistic or causal models, and a good distance between two graphs should reflect how close they are in terms of the models they represent. That is, we seek a semantic rather than structural definition of the distance. Despite these challenges, several distances have been proposed in the literature (Tsamardinos, Brown and Aliferis, 2006; Peters and Bühlmann, 2015; Viinikka, Eggeling and Koivisto, 2018), which we divide into two categories — model-agnostic and task-oriented distances.
@@
Arguably, the most popular distance is the structural Hamming distance (SHD) (Tsamardinos, Brown and Aliferis, 2006). It counts the number of edges that differ between two graphs with various proposals for how to weigh different types of edge mismatches (Perrier, Imoto and Miyano, 2008). Irrespective of the concrete weighing the SHD is computationally cheap, flexible with respect to the graph class, and guaranteed to define a metric. It is, however, model-agnostic — it disregards the specific models the graphs represent and is solely defined on the graphs themselves. This results in undesirable behaviors. First, for a pair of graphs, the SHD provides a fixed answer regardless of how the graphs are interpreted, whether probabilistically or causally. Second, the SHD does not depend on the set of graphs in consideration — it is agnostic to constraints posed by the graph type and domain specific considerations (e.g., one may wish to consider acyclic graphs). Instead the SHD treats each edge individually, which ignores that not all edge configurations may give a valid graph in the graph class under consideration. We illustrate these issues with an example.
@@
Consider the four CPDAGs listed in Fig.˜1. Each graph represents a set of distributions over $n+1$ random variables that obey the probabilistic constraints it encodes, where a missing edge between two variables signifies the conditional independence between them given another set of variables. Based on the edges (ignoring the difference between directed and undirected edges for now), we can see $\mathcal{G}_{1}\succ\mathcal{G}_{2}\succ\mathcal{G}_{3}\succ\mathcal{G}_{4}$ in terms of the models they represent. Table˜1 shows the distance of graphs $\mathcal{G}_{2},\mathcal{G}_{3},\mathcal{G}_{4}$ relative to $\mathcal{G}_{1}$: the first two columns list the two versions of the SHD (see Definition˜7) and the third column is the distance we propose in this paper. Our distance reflects what we expect logically: relative to $\mathcal{G}_{1}$, the graph $\mathcal{G}_{2}$ is the closest, followed by $\mathcal{G}_{3}$, and lastly $\mathcal{G}_{4}$. In contrast, the SHD incorrectly deems $\mathcal{G}_{3}$ to be the closest (can be much closer than $\mathcal{G}_{2}$ and $\mathcal{G}_{4}$) and fails to capture the fact that $\mathcal{G}_{2}$ is strictly closer than $\mathcal{G}_{4}$. This is due to the fact that the SHD wrongly assumes that undirecting any edge in $\mathcal{G}_{2}$ leads to a valid CPDAG that lies in between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, thereby overestimating the distance between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
Moreover, our distance agrees with the Bayesian Information Criterion (BIC) listed in the last column: we generate data under $\mathcal{G}_{1}$, specifically from a Gaussian structural equation model over $11$ variables (with unit edge weights and noise variances), and draw 100 datasets each with sample size 1000; we then fit models and estimate $\operatorname{\mathbb{E}}[\mathrm{BIC}(\hat{\mathcal{G}}_{i})-\mathrm{BIC}(\hat{\mathcal{G}}_{1})]$ (a lower score means a better model). Meanwhile, the discordance between BIC and SHD indicates that using SHD to benchmark score-based causal discovery algorithms (see Drton and Maathuis, 2017, §4) that rely on BIC or the like, such as the well-known greedy equivalence search algorithm (Chickering, 2002), can be misleading.
Table 1: Distances and BIC ($\pm$ standard error) relative to $\mathcal{G}_{1}$ for graphs in Fig.˜1 (assuming $n$ is even)
Graph $\operatorname{\mathrm{SHD}}_{1}$ $\operatorname{\mathrm{SHD}}_{2}$ Model-oriented distance BIC
$\mathcal{G}_{2}$ $n+1$ $n+2$ 11 $690\pm 5$
$\mathcal{G}_{3}$ 22 44 22 $1084\pm 6$
$\mathcal{G}_{4}$ $n/2+1$ $n+2$ $n/2+1$ $10159\pm 20$
@@
To address the limitation of model-agnostic distances, Peters and Bühlmann (2015) proposed an alternative, task-oriented approach, where the distance between two graphs is measured according to a concrete task they serve. For causal graphs, the central task is effect identification, i.e., expressing a causal effect as a functional of the observed distribution valid under any system compatible with the graph. Motivated by this, Peters and Bühlmann (2015) introduced the Structural Intervention Distance (SID), which counts the number of causal effects identified in one graph that would be invalid if the true data generating process corresponded to the other. While Henckel, Würtzen and Weichwald (2024) generalized this approach into a broader class of adjustment identification distances, the SID is not a metric: it allows zero distance between distinct graphs, lacks symmetry, and violates the triangle inequality. We show in Section˜H.1 that even a symmetrized version of SID fails the triangle inequality, limiting the utility of such task-oriented measures.
@@
Figure 1: CPDAGs $\mathcal{G}_{1}\succ\mathcal{G}_{2}\succ\mathcal{G}_{3}\succ\mathcal{G}_{4}$ in terms of the models they represent.
1.1 Main contributions and organization of the paper
We introduce a model-oriented framework that defines distance between graphs through the models they represent. Given a class of graphs of the same type (e.g., the collection of CPDAGs over five vertices), our approach begins by interpreting a probabilistic or causal graph as a statistical model, namely a set of distributions over a common collection of random variables — factual or counterfactual — that obey the constraints encoded by the graph. The notation and background for these models are described in Section˜2. We demonstrate in Section˜3 that this viewpoint naturally organizes graphs into a partially ordered set (poset) defined by model inclusion: for graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$, we say $\mathcal{G}\preceq\mathcal{G}^{\prime}$ if the model represented by $\mathcal{G}$ is contained by the model represented by $\mathcal{G}^{\prime}$. We leverage the poset’s intrinsic notion of neighborhood, called covering relations, to define the model-oriented distance as the length of a shortest path between two graphs when moving only through neighbors. By construction, this distance is a metric applicable to any class of graphs satisfying some mild assumptions, regardless of its edge types, constraints or semantics.
We note that the notion of neighborhood is also central to score-based causal discovery, where an algorithm traverses neighborhoods to search for a score-minimizing model; see, e.g., Chickering (2002); Hauser and Bühlmann (2012); Ng et al. (2024). However, to the best of our knowledge, this is the first work to leverage the structural properties of the underlying model space to formally define and analyze a distance metric between graphs.
To facilitate distance computation, in Section˜4 we analyze the structural properties of posets underlying four graph types, ordered by increasing complexity: probabilistic UGs, causal DAGs, probabilistic CPDAGs and causal MPDAGs. A key ingredient of this analysis is the graphical characterization of model containment and, where feasible, of neighborhood. In cases where characterizing the general neighborhood proves intractable, we restrict our study to the subclass of polytrees. We show that for CPDAGs and MPDAGs, this model-oriented distance can be substantially different from the structural Hamming distance. In Section˜5, we develop a generic A* algorithm to compute our distance. The algorithm is then specialized to CPDAGs and polytree MPDAGs, for which some numerical results are also presented. We conclude the paper with a brief discussion in Section˜6.
@@ 2 Background
2.1 Notation
We generally use Roman capital letters (e.g., $\mathrm{V},\mathrm{B}$) to denote a set and lowercase letters (e.g., $v_{1},b_{2}$) to denote its elements. For two sets $\mathrm{A}$ and $\mathrm{B}$, we use $\mathrm{A}\triangle\mathrm{B}:=(\mathrm{A}\setminus\mathrm{B})\cup(\mathrm{B}\setminus\mathrm{A})$ to denote their symmetric difference. We use calligraphic letters (e.g., $\mathcal{G},\mathcal{D}$) for graphs and italic capital letters (e.g., $X,Y$) for random variables.
@@ 2.2 Graphical preliminaries
We consider graphs with vertices representing random variables and edges between them representing probabilistic or causal dependencies between these variables. Basic graphical concepts are briefly described below.
2.2.1 Vertex, edge, subgraph, cycle, PDAG
We use letter $\mathrm{V}$ to denote a set of vertices, which we assume to be finite throughout; this implies that class over $\mathrm{V}$ is also finite. For convenience, $\mathrm{V}$ can be identified with $\{1,\dots,|\mathrm{V}|\}$. We write a graph over vertex set $\mathrm{V}$ with $k\geq 1$ types of edges (e.g., undirected ‘-’ and directed $\rightarrow$’) as $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$. Depending on the edge type $i$, edge set $\mathrm{E}^{i}$ is either a set of unordered tuples ($(1,2)=(2,1)$) or a set of ordered tuples ($(1,2)\neq(2,1)$). In either case, we forbid self-loops and let $\mathrm{E}^{i}\subseteq(\mathrm{V}\times\mathrm{V})\setminus\cup_{i\in\mathrm{V}}\{(i,i)\}$ for $i=1,\dots,k$. A graph $\mathcal{G}_{1}=(\mathrm{V},\mathrm{E}_{1}^{1},\dots,\mathrm{E}_{1}^{k})$ is said to be a subgraph of $\mathcal{G}_{2}=(\mathrm{V},\mathrm{E}_{2}^{1},\dots,\mathrm{E}_{2}^{k})$ if $\mathrm{E}_{1}^{i}\subseteq\mathrm{E}_{2}^{i}$ for each edge type $i$. Given a graph $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{k})$ and a subset of vertices $\mathrm{V}^{\prime}\subseteq\mathrm{V}$, the induced subgraph of $\mathrm{V}^{\prime}$ is $\mathcal{G}_{\mathrm{V}^{\prime}}:=(\mathrm{V}^{\prime},{\mathrm{E}^{1}}^{\prime},\dots,{\mathrm{E}^{k}}^{\prime})$ where ${\mathrm{E}^{i}}^{\prime}:=\mathrm{E}^{i}\cap(\mathrm{V}^{\prime}\times\mathrm{V}^{\prime})$ for $i=1,\dots,k$. A graph is simple if between every two vertices there is at most one edge (of any type). Throughout, we only consider simple graphs. A cycle is a sequence of vertices $v_{1},v_{2},\dots,v_{k}$ ($k\geq 3$) such that there is an edge between every two adjacent vertices, $v_{k}=v_{1}$ and $v_{1},\dots,v_{k-1}$ are all distinct. A directed cycle is a cycle of the form $v_{1}\rightarrow v_{2}\rightarrow\dots\rightarrow v_{k}$. In contrast, a partially directed cycle is a cycle where the edges can be directed or undirected such that (i) at least one edge is directed and (ii) all the directed edges in the cycle point in the same direction. A graph is chordal if every cycle of length 4 or more has a chord, i.e., an edge connecting two non-consecutive vertices. A partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.